# Can LLMs Reconcile Knowledge Conflicts in Counterfactual Reasoning?

### v2

@@ 2 Related Works
Multi-hop question answering has been examined in a variety of prior works as a measure of knowledge manipulation ability. (Yang et al., 2018) introduces a benchmark for measuring multi-hop question-answering abilities for context-based question-answering. Allen-Zhu & Li (2024) examines knowledge manipulation in controlled settings, finding that LLMs can struggle to solve one-step knowledge manipulation tasks without COT or extensive fine-tuning. Wang et al. (2024a), on the other hand, shows that transformers can implicitly solve multi-hop queries in the grokking regime on a synthetic knowledge-graph task. Abramov et al. (2025) extends these findings to more real settings by augmenting Wikipedia data. Yang et al. (2024b) study the reliability of latent multi-hop reasoning on real-world knowledge, finding that certain types of relations are more conducive to multi-hop reasoning than others. Biran et al. (2024) studies the mechanistic implementation of multi-hop question-answering in LLMs, finding that it tends to arise when intermediate entities can be resolved in earlier layers of the model. . Unlike prior works, which primarily focus on multi-hop reasoning involving only parametric knowledge, we study cases in which an LLM may need to combine parametric and knowledge from its context.
@@
Hybrid techniques, especially retrieval-augmented generation (RAG), attempt to merge parametric memory with retrieved information. Methods like REALM (Guu et al., 2020), RAG (Lewis et al., 2021), Dense Passage Retrieval (DPR) (Karpukhin et al., 2020), and FiD (Izacard & Grave, 2020) condition generation on external evidence, while advanced strategies such as AdaCAD (Wang et al., 2024b) and CD2 (Jin et al., 2024) offer finer-grained balancing of these knowledge sources. Goyal et al. (2025) studies the dynamics of finetuning models for context reliance, finding that instruction tuning can often worsen context reliance. Although these methods enhance the incorporation of external factual information, they are generally not designed for the distinct challenges of counterfactual reasoning. Specifically, they do not enable LLMs to accept a hypothetical premise contradicting parametric facts, then selectively retrieve relevant parametric knowledge, and subsequently perform complex multi-hop reasoning based on this newly integrated understanding.
@@
Counterfactual reasoning is increasingly vital in Natural Language Processing (NLP), especially with Large Language Models (LLMs), whose causal reasoning capabilities are an active research area (Liu et al., 2024b). While some propose LLMs as a new frontier for textual causal discovery (Kiciman et al., 2024), critical views suggest they might be "causal parrots" merely echoing training data (Zečević et al., 2023). Benchmarking efforts consistently reveal LLM limitations: CLadder assesses formal causal reasoning (Jin et al., 2023), QRData highlights struggles with data-based causal tasks (Liu et al., 2024a), and CounterBench shows poor performance on formal counterfactual inference (Chen et al., 2025). Studies like Yamin et al. (2024) further detail these issues, identifying LLM failure modes such as over-reliance on parametric knowledge or narrative shortcuts. Despite these struggles, research actively explores using LLMs for specific NLP counterfactual tasks, like generating faithful explanations (Gat et al., 2024) or alternative textual outputs via counterfactual token generation (Chatzi et al., 2024). These endeavors underscore a critical gap: a deep understanding of how and why LLMs falter when integrating their vast parametric knowledge with novel, in-context counterfactual premises, particularly in multi-hop reasoning.

### v4

@@ 2 Related Works
Multi-hop question answering has been examined in a variety of prior works as a measure of knowledge manipulation ability. (Yang et al., 2018) introduces a benchmark for measuring multi-hop question-answering abilities for context-based question-answering. Allen-Zhu and Li (2024) examines knowledge manipulation in controlled settings, finding that LLMs can struggle to solve one-step knowledge manipulation tasks without COT or extensive fine-tuning. Wang et al. (2024a), on the other hand, shows that transformers can implicitly solve multi-hop queries in the grokking regime on a synthetic knowledge-graph task. Abramov et al. (2025) extends these findings to more real settings by augmenting Wikipedia data. Yang et al. (2024b) study the reliability of latent multi-hop reasoning on real-world knowledge, finding that certain types of relations are more conducive to multi-hop reasoning than others. Biran et al. (2024) studies the mechanistic implementation of multi-hop question-answering in LLMs, finding that it tends to arise when intermediate entities can be resolved in earlier layers of the model. . Unlike prior works, which primarily focus on multi-hop reasoning involving only parametric knowledge, we study cases in which an LLM may need to combine parametric and knowledge from its context.
@@
Hybrid techniques, especially retrieval-augmented generation (RAG), attempt to merge parametric memory with retrieved information. Methods like REALM (Guu et al., 2020), RAG (Lewis et al., 2021), Dense Passage Retrieval (DPR) (Karpukhin et al., 2020), and FiD (Izacard and Grave, 2020) condition generation on external evidence, while advanced strategies such as AdaCAD (Wang et al., 2024b) and CD2 (Jin et al., 2024) offer finer-grained balancing of these knowledge sources. Goyal et al. (2025) studies the dynamics of finetuning models for context reliance, finding that instruction tuning can often worsen context reliance. Although these methods enhance the incorporation of external factual information, they are generally not designed for the distinct challenges of counterfactual reasoning. Specifically, they do not enable LLMs to accept a hypothetical premise contradicting parametric facts, then selectively retrieve relevant parametric knowledge, and subsequently perform complex multi-hop reasoning based on this newly integrated understanding.
@@
Counterfactual reasoning is increasingly vital in Natural Language Processing (NLP), especially with Large Language Models (LLMs), whose causal reasoning capabilities are an active research area (Liu et al., 2025). While some propose LLMs as a new frontier for textual causal discovery (Kiciman et al., 2024), critical views suggest they might be "causal parrots" merely echoing training data (Zečević et al., 2023). Benchmarking efforts consistently reveal LLM limitations: CLadder assesses formal causal reasoning (Jin et al., 2023), QRData highlights struggles with data-based causal tasks (Liu et al., 2024), and CounterBench shows poor performance on formal counterfactual inference (Chen et al., 2025). Studies like Yamin et al. (2025) further detail these issues, identifying LLM failure modes such as over-reliance on parametric knowledge or narrative shortcuts. Despite these struggles, research actively explores using LLMs for specific NLP counterfactual tasks, like generating faithful explanations (Gat et al., 2024) or alternative textual outputs via counterfactual token generation (Chatzi et al., 2024). These endeavors underscore a critical gap: a deep understanding of how and why LLMs falter when integrating their vast parametric knowledge with novel, in-context counterfactual premises, particularly in multi-hop reasoning.